OpenStudy (anonymous):

Simplify (10x-3)(4x+5)

OpenStudy (anonymous):

40x^2 +38x -15

OpenStudy (anonymous):

(10x-3)(4x+5) 10x*4x=40x^2 10x*5=50x -3*4x=-12x -3*5=-15 40x^2+50x-12x-15 40x^2+38x-15

OpenStudy (anonymous):

thanks peps

OpenStudy (anonymous):

@Twitter @KissMeImScooty

OpenStudy (anonymous):

????